In a Young’s double slit experiment λ = 500 nm, d = 1.0 mm and D = 1.0 m. Find the minimum distance from the central maximum for which the intensity is half of the maximum intensity.

Grade:12

1 Answers

Arun
4 years ago
Given that, D = 1 m, d = 1 mm = 10–3 m, λ= 500 nm = 5 x10–7 m
 
y = lambda D/ 4d
 
y = 5 * 10^-7 * 1/ 4 * 10^-3
 
 = 1.25 * 10^-4 m
 
Hope ot helps